# Tunable Entanglement in Cavity-Magnon Optomechanics

### v2

@@ Abstract
Magnons Prabhakar and Stancil (2009); Kittel (1958), quanta of collective spin-wave excitations in magnetically ordered materials like yttrium iron garnet (YIG), have garnered considerable interest in quantum physics and condensed matter physics Rameshti et al. (2022); Yuan et al. (2022); Zheng et al. (2023); Li et al. (2020); Lachance-Quirion et al. (2019); Zuo et al. (2023). With unique properties such as high spin density and low energy loss Huebl et al. (2013); Tabuchi et al. (2014); Zhang et al. (2014), diverse novel phenomena including strong Huebl et al. (2013); Tabuchi et al. (2014); Zhang et al. (2014); Goryachev et al. (2014) and ultrastrong coupling Ghirri et al. (2023); Wang et al. (2024); Kostylev et al. (2016), qubit-magnon coupling Tabuchi et al. (2015), magnon dark modes and gradient memory Zhang et al. (2015), magnon Kerr effect Wang et al. (2016), spin current manipulation Bai et al. (2017), exceptional points Zhang et al. (2017); Zhang and You (2019); Harder et al. (2017); Liu et al. (2019), axion detection Mitridate et al. (2020); Ikeda et al. (2022); Barbieri et al. (1989), superadiant phase transition Liu et al. (2023), dissipative coupling Wang and Hu (2020); Harder et al. (2021), and strong spin coupling Xiong et al. (2022a, 2023); Tian et al. (2023); Neuman et al. (2020); Hei et al. (2023); Fukami et al. (2021); Hetényi et al. (2022). Recently, special attention has been paid to macroscopic quantum entanglement associated with magnons Li et al. (2018); Zhang et al. (2019); Chen et al. (2023, 2024); Yuan et al. (2020); Li and Zhu (2019); Sun et al. (2021); Ren et al. (2022); Lai et al. (2022). This is because it is crucial for understanding the classical-quantum boundary Haroche (1998) and can be regarded as an important resource for quantum information science Bouwmeester et al. (2000). However, to produce such macroscopic entanglement, nonlinear effects are always prerequired Adesso and Illuminati (2007a). The extensively explored cavity optomechanics Aspelmeyer et al. (2014); Kippenberg and Vahala (2007, 2008); Metcalfe (2014); Liu et al. (2013a); Chang and Zhang (2022); Barzanjeh et al. (2022); Xiong et al. (2015), hybridized by photons in a cavity mode weakly coupled to phonons in a mechanical resonator through radiation pressure Vitali et al. (2007), can provide the natural nonlinearity for investigating optomechanical entanglement Sarma et al. (2021) and other intriguing effects such as sensing Li et al. (2021); Liu et al. (2021); Schliesser et al. (2008), ground-state cooling Marquardt et al. (2008); Otterstrom et al. (2018); Gan et al. (2019); Park and Wang (2009); Xu et al. (2019); Liu et al. (2014); Huang et al. (2022); Genes et al. (2008); Zhang et al. (2020); Nunnenkamp et al. (2010), squeezed light generation Zhang et al. (2020); Nunnenkamp et al. (2010); Purdy et al. (2013); Aggarwal et al. (2020); Kronwald et al. (2014); Lü et al. (2015a); Liao and Law (2011); Lü et al. (2015b), nonreciprocity Mercier de Lépinay et al. (2020); Shen et al. (2023); Xu et al. (2020); Eshaqi-Sani et al. (2022); Hafezi and Rabl (2012); Peterson et al. (2017); Miri et al. (2017); Jiao et al. (2020); Xu et al. (2018), optomechanically induced transparency Yan (2020); Lü et al. (2018); Kronwald and Marquardt (2013); Dong et al. (2013); Zhang et al. (2018); Lai et al. (2020); Koppenhöfer et al. (2023); Karuza et al. (2013); Liu et al. (2013b); Weis et al. (2010), coupling enhancement Xiong et al. (2021a); Lü et al. (2013); Chen et al. (2021); Shen et al. (2016); Fang et al. (2017), multistability Xiong et al. (2016); Kyriienko et al. (2014); Sete and Eleuch (2012); Ghobadi et al. (2011), and high-order exceptional points Xiong et al. (2022b, 2021b); Jing et al. (2017). This indicates that combination of magnonics and optomechanics is a potential path to investigate macroscopic magnon entanglement.
@@
Owing to the strong photon-magnon coupling in planar Huebl et al. (2013) and 3D cavities Tabuchi et al. (2014); Zhang et al. (2014); Goryachev et al. (2014), hybridizing magnons with optomechanics to form cavity-magnon optomechanics becomes feasible Shen et al. (2022); Chen et al. (2024); Gao et al. (2017); Chen et al. (2023); Xiong et al. (2023) for utilizing their individual advantages. In this work, we propose to generate diverse bipartite and tripartite entanglement in hybrid cavity-magnon optomechanics. Compared to the standard cavity optomechanics, the optomechanical entanglement can be adjusted flexibly by the introduced magnons in the Kittle mode of a single sphere. Specifically, the optimal optomechanical entanglement can be predicted by tuning the detuning of the Kittle mode from the driving field to be resonant with the mechanical mode (i.e., $\Delta_{m}=\omega_{b}$). But when we tune this detuning to be anti-resonant with the mechanical mode (i.e., $\Delta_{m}=-\omega_{b}$), the optomechanical entanglement is weakest, while the magnon-photon entanglement, the magnon-phonon, as well as the magnon-photon-phonon entanglement are approximately optimal. These entanglement can be further enhanced by engineering the coupling between magnons and photons. In addition, the tunable magnon-magnon entanglement and the tripartite entanglement including the magnon-magnon-photon entanglement and the magnon-magnon-phonon entanglement can also be generated in the proposed cavity-magnon optomechanics with two spheres. Tuning one Kittle mode to be resonant with the mechanical mode (i.e., $\Delta_{1}=\omega_{b}$), while the other to be anti-resonant with the mechanical mode (i.e., $\Delta_{2}=-\omega_{b}$), or vice versa, the optimal magnon-magnon entanglement is predicted, and its value approximately equals to half the optomechanical entanglement. But to have optimal magnon-magnon-photon and -phonon entanglement, two Kittle modes are needed to be resonant (i.e., $\Delta_{1}=\Delta_{2}$). We further show that the optomechanical entanglement can be protected or restored by using bad magnons with large decay rate, but the magnon-magnon entanglement as well as the magnon-magnon-photon and -phonon entanglement are severly reduced. Besides, the bath temperature effect on these entanglement are also investigated. We find that for the weaker magnon-photon coupling (i.e., $g_{m}<G$), the stronger optomechanical entanglement than that of the magnon-magnon entanglement can be obtained within the survival temperature, but the situation is reversed when the magnon-photon coupling becomes comparable with the optomechanical coupling (i.e., $g_{m}\approx G$). Moreover, the survival temperature of the magnon-magnon entanglement is more robust against the magnon-photon coupling than that of the optomechanical entanglement. For the tripartite entanglement, we show that the magnon-magnon-phonon entanglement is stronger than the magnon-magno-photon entanglement within their survival temperature, owing to the mechanical cooling. All achieved entanglement exhibit robust against the bath temperature. Our proposal opens an alternative path to study tunable macroscopic quantum effect in hybrid cavity-magnon optomechanics.
@@
This rest paper is organized as follows: In Sec. II, the model is described and the entanglement quantification is given. Then in Sec. III, we study optomechanical entanglement mediated by magnons in the proposed cavity-magnon optomechanics, where only a single sphere is considered. In Sec. IV, tunable magnon-magnon entanglement induced by the optomechanical interface is investigated via examing the effect of system parameters. In Sec. V, we further study tripartite entanglement including magnon-magnon-photon and magnon-magnon-phonon entanglement. Finally, a conclusion is given in Sec. VI.
@@ II.1 Model
Figure 1: (a) Schematic diagram of the proposed cavity-magnon optomechanics. It consists of two YIG spheres placed in biased magnetic field $B_{0}$, a driven cavity and a mechanical resonator. The cavity mode is strongly coupled to the Kittle modes of two YIG spheres, and weakly coupled to the mechanical mode. (b) The effective coupling configuration. $G$ is the linearized optomechanical coupling strength, and $g_{1(2)}$ is the tunable coupling strength between the cavity mode and the Kittle mode of the sphere $1~{}(2)$. $a,b,m$ denote the cavity mode, the mechanical mode, and the Kittle mode, respectively.
@@
where $H_{\rm OM}=\Delta_{a}a^{\dagger}a+\frac{\omega_{b}}{2}\left(q^{2}+p^{2}\right)+g_{0}a^{\dagger}aq$ denotes the Hamiltonia of the standard cavity optomechanics, describing the interaction between the cavity mode and the mechanical mode via the radiation pressure, with the single-photon optomechanical coupling strength $g_{0}$. $\Delta_{a}=\omega_{a}-\omega_{0}$, with $\omega_{a}$ being the frequency of the cavity mode, is the frequency detuning of the cavity mode from the driving field, and $\omega_{b}$ is the frequency of the mechanical mode. The Hamiltonian $H_{\rm CM}^{(j)}=\Delta_{j}m^{\dagger}_{j}m_{j}+g_{j}(m_{j}a^{\dagger}+m_{j}^{\dagger}a)$ charaterizes the interaction between the cavity mode and the Kittle mode of the $j$th sphere. $\Delta_{j}=\omega_{j}-\omega_{0}$ is the frequency detuning of the Kittle mode of the $j$th sphere from the driving field. The creation (annihilation) operators the cavity mode and the Kittle mode of the $j$th sphere are denoted by $a^{\dagger}$ ($a$) and $m_{j}^{\dagger}$ ($m_{j}$), respectively. $q$ and $p$ are dimensionless position and momentum quadratures of the mechanical mode.
@@ II.2 Quantum Langevin equation
$\dot{p}=$ $-\omega_{b}q-g_{0}a^{\dagger}a-\gamma_{b}p+\xi,$
$\dot{m_{j}}=$ $-(\kappa_{j}+i\Delta_{j})m_{j}-ig_{j}a+\sqrt{2\kappa_{j}}\zeta_{j},$ (2)
$\dot{a}=$ $-(\kappa_{a}+i{\Delta}_{a})a-ig_{0}aq-ig_{j}m_{j}+\Omega+\sqrt{2\kappa_{a}}\mu,$
where $\kappa_{a}$, $\kappa_{j}$, and $\gamma_{b}$ are the decay rates of the cavity mode, the Kittle mode of the $j$th sphere, and the mechanical mode, respectively. For simplicity, we assume that the Kittle modes of two spheres have the identical decay rate $\kappa_{1}=\kappa_{2}=\kappa_{m}$ below. $\mu$, $\zeta_{j}$ and $\xi$ are the vaccum input noise operators for the cavity mode, the mechanical mode, and the Kittle mode of the $j$th sphere, having the zero mean values $\langle\mu\rangle=\langle\zeta_{j}\rangle=\langle\xi\rangle=0$. Under the Markovian approximation, two-time correlation functions of these input noise operators are
@@
$\langle\xi(t)\xi(t^{\prime})+\xi(t^{\prime})\xi(t)\rangle\simeq$ $2\gamma_{b}(2\bar{n}_{b}+1)\delta(t-t^{\prime}),$
where $\bar{n}_{\sigma}=\left[\exp\left(\hbar\omega_{\sigma}/k_{B}T\right)-1\right]^{-1}$ ($\sigma=a,b,j$) is the mean thermal excitation number, with $k_{B}$ being the Boltzmann constant and $T$ the bath temperature.
@@
where $\tilde{\Delta}_{a}={\Delta}_{a}+g_{0}\langle q\rangle$ is the effective frequency detuning, induced by the displacement of the mechanical mode, and $G=\sqrt{2}g_{0}\langle a\rangle$ is the effectively enhanced optomechanical coupling strength between the cavity mode and the mechanical mode by the factor $\langle a\rangle$. The steady-state values $\langle O\rangle$ in Eq. (II.2) are given by $\dot{\langle O\rangle}=0$. Specifically, $\langle p\rangle=0$, $\langle q\rangle=-{g_{0}|\langle a\rangle|^{2}}/{\omega_{b}}$, $\langle m_{j}\rangle\approx-{g_{j}\langle a\rangle}/{\Delta_{j}}$, and $\langle a\rangle\approx{i\Omega}/{(g_{j}^{2}/\Delta_{j}-\tilde{\Delta}_{a})}$, where $\left|\Delta_{j}\right|,\left|\tilde{\Delta}_{a}\right|\gg\kappa_{j},\kappa_{a}$ are taken. Experimentally, these conditions can be easily achieved by tuning the frequency of the driving field. Also, the value of $\langle a\rangle$ can be real via tuning the initial phase of the strong driving field. Hereafter, we assume $\langle a\rangle$ to be real for simplicity. Thus, the effective Hamiltonian of the linearized hybrid system can be written as

### v4

# Hybrid Cavity-Magnon Optomechanics: Tailoring Bipartite and Tripartite Macroscopic Entanglement

@@ Abstract
Magnons Prabhakar and Stancil (2009); Kittel (1958), quanta of collective spin-wave excitations in magnetically ordered materials like yttrium iron garnet (YIG), have garnered considerable interest in quantum physics and condensed matter physics Rameshti et al. (2022); Yuan et al. (2022); Zheng et al. (2023); Li et al. (2020); Lachance-Quirion et al. (2019); Zuo et al. (2024). With unique properties such as high spin density and low energy loss, diverse novel phenomena including strong Huebl et al. (2013); Tabuchi et al. (2014); Zhang et al. (2014); Goryachev et al. (2014) and ultrastrong coupling Ghirri et al. (2023); Wang et al. (2024); Kostylev et al. (2016), qubit-magnon coupling Tabuchi et al. (2015), magnon dark modes and gradient memory Zhang et al. (2015), magnon Kerr effect Wang et al. (2016), spin current manipulation Bai et al. (2017), exceptional points Zhang et al. (2017); Zhang and You (2019); Harder et al. (2017); Liu et al. (2019), axion detection Mitridate et al. (2020); Ikeda et al. (2022); Barbieri et al. (1989), superadiant phase transition Liu et al. (2023), dissipative coupling Wang and Hu (2020); Harder et al. (2021), and strong spin coupling Xiong et al. (2022a, 2023); Peng et al. (2025); Neuman et al. (2020); Hei et al. (2023); Fukami et al. (2021); Hetényi et al. (2022); Chen et al. (2025). Very recently, a single magnon manipulation Xu et al. (2023, 2024) and quantum magnon squeezing You et al. have been demonstrated. This greatly promotes the further development of magnonics. In magnonics, how to prepare macroscopic entanglement is a long outstanding problem Li et al. (2018); Zhang et al. (2019); Chen et al. (2023, 2024); Yuan et al. (2020); Li and Zhu (2019); Sun et al. (2021); Ren et al. (2022); Lai et al. (2022); Liu et al. (2025a); Yu et al. (2020); Yang et al. (2021); Tan and Li (2021); Zhou et al. (2022); Noura et al. (2024). This is because it is crucial for understanding the classical-quantum boundary Haroche (1998) and can be regarded as an important resource for quantum information science Bouwmeester et al. (2000). However, to produce such macroscopic entanglement, nonlinear effects are always prerequired Adesso and Illuminati (2007a). The extensively explored cavity optomechanics Aspelmeyer et al. (2014); Kippenberg and Vahala (2007, 2008); Metcalfe (2014); Liu et al. (2013a); Chang and Zhang (2022); Barzanjeh et al. (2022); Xiong et al. (2015); Amazioug et al. (2020), hybridized by photons in a cavity mode weakly coupled to phonons in a mechanical resonator through radiation pressure Vitali et al. (2007), can provide the natural nonlinearity for investigating optomechanical entanglement Sarma et al. (2021) and other intriguing effects such as sensing Li et al. (2021); Liu et al. (2021); Schliesser et al. (2008), ground-state cooling Marquardt et al. (2008); Otterstrom et al. (2018); Gan et al. (2019); Park and Wang (2009); Xu et al. (2019); Liu et al. (2014); Huang et al. (2022); Genes et al. (2008); Zhang et al. (2020); Nunnenkamp et al. (2010); Fan et al. (2025), squeezed light generation Purdy et al. (2013); Aggarwal et al. (2020); Kronwald et al. (2014); Lü et al. (2015a); Liao and Law (2011); Lü et al. (2015b), nonreciprocity Mercier de Lépinay et al. (2020); Shen et al. (2023); Xu et al. (2020); Eshaqi-Sani et al. (2022); Hafezi and Rabl (2012); Peterson et al. (2017); Miri et al. (2017); Jiao et al. (2020); Xu et al. (2018), optomechanically induced transparency Yan (2020); Lü et al. (2018); Kronwald and Marquardt (2013); Dong et al. (2013); Zhang et al. (2018); Lai et al. (2020); Koppenhöfer et al. (2023); Karuza et al. (2013); Liu et al. (2013b); Weis et al. (2010), coupling enhancement Xiong et al. (2021a); Lü et al. (2013); Chen et al. (2021); Shen et al. (2016); Fang et al. (2017), multistability Xiong et al. (2016); Kyriienko et al. (2014); Sete and Eleuch (2012); Ghobadi et al. (2011), and high-order exceptional points Xiong et al. (2022b, 2021b); Jing et al. (2017). This indicates that combination of magnonics and optomechanics is a potential path to investigate macroscopic magnon entanglement.
@@
Owing to the strong photon-magnon coupling in planar Huebl et al. (2013) and 3D cavities Tabuchi et al. (2014); Zhang et al. (2014); Goryachev et al. (2014), hybridizing magnons with optomechanics to form cavity-magnon optomechanics becomes feasible Shen et al. (2022); Chen et al. (2024); Gao et al. (2017); Chen et al. (2023); Xiong et al. (2023); Sohail et al. (2025); Mathkoor et al. (2025); Hidki et al. (2024); Ahmed et al. (2025) for utilizing their individual advantages. In this work, we propose to generate diverse bipartite and tripartite entanglement in hybrid cavity-magnon optomechanics. Compared to the standard cavity optomechanics, the optomechanical entanglement can be adjusted flexibly by the introduced magnons in the Kittle mode of a single sphere. Specifically, the optimal optomechanical entanglement can be predicted by tuning the detuning of the Kittle mode from the driving field to be resonant with the mechanical mode (i.e., $\Delta_{m}=\omega_{b}$). But when we tune this detuning to be anti-resonant with the mechanical mode (i.e., $\Delta_{m}=-\omega_{b}$), the optomechanical entanglement is weakest, while the magnon-photon entanglement, the magnon-phonon, as well as the magnon-photon-phonon entanglement are approximately optimal. These entanglement can be further enhanced by engineering the coupling between magnons and photons. In addition, the tunable magnon-magnon entanglement and the tripartite entanglement including the magnon-magnon-photon entanglement and the magnon-magnon-phonon entanglement can also be generated in the proposed cavity-magnon optomechanics with two spheres. Tuning one Kittle mode to be resonant with the mechanical mode (i.e., $\Delta_{1}=\omega_{b}$), while the other to be anti-resonant with the mechanical mode (i.e., $\Delta_{2}=-\omega_{b}$), or vice versa, the optimal magnon-magnon entanglement is predicted, and its value approximately equals to half the optomechanical entanglement. But to have optimal magnon-magnon-photon and -phonon entanglement, two Kittle modes are needed to be resonant (i.e., $\Delta_{1}=\Delta_{2}$). We further show that the optomechanical entanglement can be protected or restored by using bad magnons with large decay rate, but the magnon-magnon entanglement as well as the magnon-magnon-photon and -phonon entanglement are severly reduced. Besides, the bath temperature effect on these entanglement are also investigated. We find that for the weaker magnon-photon coupling (i.e., $g_{m}<G$), the stronger optomechanical entanglement than that of the magnon-magnon entanglement can be obtained within the survival temperature, but the situation is reversed when the magnon-photon coupling becomes comparable with the optomechanical coupling (i.e., $g_{m}\approx G$). Moreover, the survival temperature of the magnon-magnon entanglement is more robust against the magnon-photon coupling than that of the optomechanical entanglement. For the tripartite entanglement, we show that the magnon-magnon-phonon entanglement is stronger than the magnon-magnon-photon entanglement within their survival temperature, owing to the mechanical cooling. All achieved entanglement exhibit robustness against the bath temperature. Our proposal opens an alternative path to study tunable macroscopic quantum effect in hybrid cavity-magnon optomechanics.
@@
The rest of the paper is organized as follows: In Sec. II, the model is described and the entanglement quantification is given. Then in Sec. III, we study optomechanical entanglement mediated by magnons in the proposed cavity-magnon optomechanics, where only a single sphere is considered. In Sec. IV, tunable magnon-magnon entanglement induced by the optomechanical interface is investigated via examing the effect of system parameters. In Sec. V, we further study tripartite entanglement including magnon-magnon-photon and magnon-magnon-phonon entanglement. Finally, a conclusion is given in Sec. VI.
@@ II.1 Model
Figure 1: (a) Schematic diagram of the proposed cavity-magnon optomechanics. It consists of two YIG spheres placed in biased magnetic field $B_{0}$, a driven cavity and a mechanical resonator. The cavity mode is strongly coupled to the Kittle modes of two YIG spheres, and weakly coupled to the mechanical mode. (b) The effective coupling configuration. $G$ is the linearized optomechanical coupling strength, and $g_{1(2)}$ is the tunable coupling strength between the cavity mode and the Kittle mode of the YIG sphere 1(2)1~(2).
@@
where $H_{\rm OM}=\Delta_{a}a^{\dagger}a+\frac{\omega_{b}}{2}\left(q^{2}+p^{2}\right)+g_{0}a^{\dagger}aq$ denotes the Hamiltonian of the standard cavity optomechanics, describing the interaction between the cavity mode and the mechanical mode via the radiation pressure, with the single-photon optomechanical coupling strength $g_{0}$. $\Delta_{a}=\omega_{a}-\omega_{0}$, with $\omega_{a}$ being the frequency of the cavity mode, is the frequency detuning of the cavity mode from the driving field, and $\omega_{b}$ is the frequency of the mechanical mode. The Hamiltonian $H_{\rm CM}^{(j)}=\Delta_{j}m^{\dagger}_{j}m_{j}+g_{j}(m_{j}a^{\dagger}+m_{j}^{\dagger}a)$ charaterizes the interaction between the cavity mode and the Kittle mode of the $j$th sphere. $\Delta_{j}=\omega_{j}-\omega_{0}$ is the frequency detuning of the Kittle mode of the $j$th sphere from the driving field. The creation (annihilation) operators the cavity mode and the Kittle mode of the $j$th sphere are denoted by $a^{\dagger}$ ($a$) and $m_{j}^{\dagger}$ ($m_{j}$), respectively. $q$ and $p$ are dimensionless position and momentum quadratures of the mechanical mode.
@@ II.2 Quantum Langevin equation
$\dot{p}=$ $-\omega_{b}q-g_{0}a^{\dagger}a-\gamma_{b}p+\xi(t),$
$\dot{m_{j}}=$ $-(\kappa_{j}+i\Delta_{j})m_{j}-ig_{j}a+\sqrt{2\kappa_{j}}\zeta_{j}(t),$ (2)
$\dot{a}=$ $-(\kappa_{a}+i{\Delta}_{a})a-ig_{0}aq-ig_{j}m_{j}+\Omega+\sqrt{2\kappa_{a}}\mu(t),$
where $\kappa_{a}$, $\gamma_{b}$, and $\kappa_{m,j}$ denote the decay rates of the cavity mode, the mechanical mode, and the Kittel mode of the $j$th YIG sphere, respectively. The corresponding vacuum input noise operators are $\mu(t)$, $\xi(t)$, and $\zeta_{j}(t)$, which satisfy zero mean values $\langle\mu(t)\rangle=\langle\xi(t)\rangle=\langle\zeta_{j}(t)\rangle=0$. Under the Markovian approximation, their two-time correlation functions are given by
@@
$\langle\xi(t)\xi(t^{\prime})+\xi(t^{\prime})\xi(t)\rangle$ $\simeq 2\gamma_{b}(2\bar{n}_{b}+1)\delta(t-t^{\prime}),$
where $\bar{n}_{\sigma}=[\exp(\hbar\omega_{\sigma}/k_{B}T)-1]^{-1}$ ($\sigma=a,b,j$) is the mean thermal excitation number, $k_{B}$ is the Boltzmann constant, and $T$ is the bath temperature.
@@
where $\tilde{\Delta}_{a}={\Delta}_{a}+g_{0}\langle q\rangle$ is the effective frequency detuning, induced by the displacement of the mechanical mode, and $G=\sqrt{2}g_{0}\langle a\rangle$ is the effectively enhanced optomechanical coupling strength between the cavity mode and the mechanical mode by the factor $\langle a\rangle$. The steady-state values $\langle O\rangle$ in Eq. (II.2) are given by $\dot{\langle O\rangle}=0$. Specifically, $\langle p\rangle=0$, $\langle q\rangle=-{g_{0}|\langle a\rangle|^{2}}/{\omega_{b}}$, $\langle m_{j}\rangle\approx-{g_{j}\langle a\rangle}/{\Delta_{j}}$, and $\langle a\rangle\approx{i\Omega}/{(g_{j}^{2}/\Delta_{j}-\tilde{\Delta}_{a})}$, where $\left|\Delta_{j}\right|,\left|\tilde{\Delta}_{a}\right|\gg\kappa_{j},\kappa_{a}$ are taken. Experimentally, these conditions can be easily achieved by tuning the frequency of the driving field Wang et al. (2016). Also, the value of $\langle a\rangle$ can be real via tuning the initial phase of the strong driving field. Hereafter, we assume $\langle a\rangle$ to be real for simplicity, directly giving rise to $G=G^{*}$. Thus, the effective Hamiltonian of the linearized hybrid system can be written as